# Investigation, Writing - Review & Editing

### v1

@@ 4.1 Implementation Details
| Ours | 97.9 | 80.2 |
| Methods | OA | mIoU | Preprocessing (s) | Inference (s) | Model Parameters (M) |
| SPT(Robert et al., 2023) | 97.5 | 79.6 | 125.3 | 2.0 | 0.8 |
| 3DLST | 97.9 | 80.2 | 76.1 | 2.36 | 16.2 |
@@ 4.3 DALES Dataset
Performance Comparison. Comparison results of different methods on the DALES dataset are shown in Table 3. We compared 3DLST with other SOTA point cloud segmentation methods including the most recent superpoint Transformer method, SPT (Robert et al., 2023). From the results, our method outperforms other benchmarked SOTA methods, with the highest mIoU (80.2) and OA (97.9). This demonstrates the excellent segmentation performance of 3DLST to aerial LiDAR data. Moreover, we also compared the inference speed of 3DLST and SPT (Robert et al., 2023) on one testing area () in detail. As shown in Table 4, 3DLST is over 2 faster than SPT in terms of the preprocessing step. This is because 3DLST does not require the time-consuming superpoint generation process. Since the size of input tokens to SPT is much smaller than ours, the model parameters of SPT are much fewer than ours. Therefore, the network inference speed of 3DLST is slightly lower than SPT. Overall, the entire inference speed of 3DLST is 1.6 faster than SPT, while achieving a slightly higher mIoU. This demonstrates that 3DLST achieves SOTA performance in both algorithm accuracy and speed. The visualization of the segmentation result comparison on the DALES dataset is shown in Fig. 4.1, which confirms the superiority of 3DLST in aerial LiDAR point cloud segmentation.
@@ 4.5 Ablation Study
| Cross-Attention-guided Upsampling | Trilinear Interpolation | 86.3 | 77.5 | 93.7 | 12.7 |
| Nearest Neighbor Interpolation | 86.0 | 77.6 | 94.3 | 12.0 | |
@@
Cross-Attention-guided Upsampling. We conducted ablation studies for the proposed CAU block to show its superiority compared with other common upsampling methods. Specifically, we replaced the CAU block with the trilinear and nearest neighbor interpolation methods respectively. As shown in Table 6 Row 4 and 5, the CAU block achieved better results in terms of all accuracy metrics, as well as with lower latency (11.7 ). The clustering relationships stored in CAM allow the CAU block to achieve an accurate and efficient process of token reconstruction.

### v2

# 3D Learnable Supertoken Transformer for LiDAR Point Cloud Scene Segmentation

@@ 4.1 Implementation Details
| Ours | 97.6 | 80.2 |
| Methods | OA | mIoU | Preprocessing (s) | Inference (s) |
| SPT(Robert et al., 2023) | 97.5 | 79.6 | 125.3 | 2.0 |
| 3DLST | 97.6 | 80.2 | 76.1 | 2.36 |
@@ 4.3 DALES Dataset
Performance Comparison. Comparison results of different methods on the DALES dataset are shown in Table 3. We compared 3DLST with other SOTA point cloud segmentation methods including the most recent superpoint Transformer method, SPT (Robert et al., 2023). From the results, our method outperforms other benchmarked SOTA methods, with the highest mIoU (80.2) and OA (97.6). This demonstrates the excellent segmentation performance of 3DLST to aerial LiDAR data. Moreover, we also compared the inference speed of 3DLST and SPT (Robert et al., 2023) on one testing area () in detail. As shown in Table 4, 3DLST is over 2 faster than SPT in terms of the preprocessing step. This is because 3DLST does not require the time-consuming superpoint generation process. Since the size of input tokens to SPT is smaller than ours (by the preprocessing step of superpoint clustering), the network inference speed of 3DLST is slightly lower than SPT. Overall, the entire inference speed of 3DLST is 1.6 faster than SPT, while achieving a slightly higher mIoU. This demonstrates that 3DLST achieves SOTA performance in both algorithm accuracy and speed. The visualization of the segmentation result comparison on the DALES dataset is shown in Fig. 4.1, which confirms the superiority of 3DLST in aerial LiDAR point cloud segmentation.